OpenStudy (anonymous):

The amount of paint needed to cover a wall is proportional to its area. The wall is rectangular and has an area of 4z^2+2z square meters. Factor this polynomial to find possible expressions for the length and width of the wall. (Assume the factors are polynomials.) If z=4 what would the length and width be?

OpenStudy (anonymous):

please help me understand this.

OpenStudy (anonymous):

if z=4, then what is the area?

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

4z^2+2z plug in 4 for z

OpenStudy (anonymous):

264 ?

OpenStudy (anonymous):

no. 4(4)^2=4*16=64 2*4=8 64+8=72

OpenStudy (anonymous):

oh, I squared it again 16^2+8, I get it

OpenStudy (anonymous):

now the question says to factor the polynomial. can you do that?

OpenStudy (anonymous):

No, how would I do that?

OpenStudy (anonymous):

ok. we need 2 different numbers which when multiplied together equals 72. so we factor 4z^2+2z by "taking out" of each term the common factors. what do you think the common factors are?

OpenStudy (anonymous):

2 and 4?

OpenStudy (anonymous):

getting close. 2 is common to each of 4z^2 and 2 z. 4 is not common to both. but there is something else that they share in common.

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

z is comon to both.

OpenStudy (anonymous):

oh, ok. both of them have z

OpenStudy (anonymous):

so "take out" 2z from both terms as follows 2z(?+?)=4z^2+2z

OpenStudy (anonymous):

(?+?)=4z^2 like this?

OpenStudy (anonymous):

like this 2z(2z+1)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so 2z is one number and 2z+1 is the other number. if z=4, what are the 2 numbers and if you multiply them together do you get 72?

OpenStudy (anonymous):

8 and 9

OpenStudy (anonymous):

yup. good job!

OpenStudy (anonymous):

thanks alot!

OpenStudy (anonymous):

you are welcome!